Grade 11Physical Chemistry

Calculate the no. of moles of electrons, protons, and neutrons present in 18 grams of Water

Nitin Manohar Mishra
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
 = Avogadro’s numberA electronsNA moleculestotal electrons in water = 2(H) + 8(O) = 101 molecule of water contains = 10 electrons1 mole water contains = 10 * NADear studentHere we will use the the concept of density of water = 1g/ccD=M/V1 g/ml = M/18 mlM= 18 gmoles of water = mass/molar mass = 18/18 = 11 mole of water contains = N